OpenStudy (anonymous):

Solve the system, using substitution. One point for correct steps, one point for correct ordered pair. x = 4y and 2x + 3y = 44

OpenStudy (kidrah69):

\[2(4y)+3y=44\] i plugged in x=4y in place of X

OpenStudy (kidrah69):

can u solve the rest?

OpenStudy (kidrah69):

Combine the like terms....whats 4y+3y ?

OpenStudy (anonymous):

no can u walk me through it

OpenStudy (kidrah69):

i just took x=4y and plugged it into the other equation nothing complicated

OpenStudy (kidrah69):

\[2(4y)+3y=44 \] 4y+3y=?

OpenStudy (anonymous):

7y

OpenStudy (kidrah69):

good! now we have \[2+7y=44\]

OpenStudy (kidrah69):

subtract 2 from both sides what are you left with?

OpenStudy (anonymous):

i got 44/9

OpenStudy (kidrah69):

thats incorrect if you subtract 44-2 you get 42 what is 42/7 ?

OpenStudy (anonymous):

6

OpenStudy (kidrah69):

good so y=6

OpenStudy (kidrah69):

now we plug that in to the other equation

OpenStudy (anonymous):

ok

OpenStudy (kidrah69):

\[x=4(6)\]

OpenStudy (anonymous):

i got 24

OpenStudy (anonymous):

is that right

OpenStudy (kidrah69):

yep x=24 y=6

OpenStudy (anonymous):

the what

OpenStudy (kidrah69):

what?

OpenStudy (anonymous):

is that all for the question

OpenStudy (kidrah69):

yeah :)

OpenStudy (anonymous):

so i got it !

OpenStudy (anonymous):

i got 2 go see u tomorrow.

OpenStudy (kidrah69):

bye :) and good work!